OpenStudy (chrisasl):

Hello, Is there any theorem for the rank of the product of two matrices A,B (both of them are nxn dimension)? I mean something like rank(AB) = rank(BA) when A,B matrices are of the same dimension nxn?

OpenStudy (anonymous):

I think you already know these, but what I can remember is that, if B is a matrix n-by-k: rank(AB) = min(rank(A), rank(B)) If B has rank n, then: rank(AB) = rank(A) and if C is a matrix k-by-m with rank m: rank(CA) = rank(A). Sorry if you already know these, LA is not my strong subject :-)

OpenStudy (chrisasl):

I've found these properties of rank at wikipedia, already! Thanks for answering, though :-)

OpenStudy (anonymous):

No problem :-)